OpenStudy (anonymous):

1) An Olympic archer is able to hit the bull’s-eye 80% of the time. Assume that each shot is independent of the others. If she shoots 6 arrows, what is the probability of each result below? a) She gets exactly 4 bull’s-eyes.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

the question is what is the probability that she gets exactly 4

OpenStudy (anonymous):

o

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

well there i cant help you sorry

OpenStudy (anonymous):

@jim_thompson5910 this person needs help and i cant help them but they have been trying for a long time can you help

jimthompson5910 (jim_thompson5910):

Use the binomial probability distribution (n C k)*(p)^(k)*(1-p)^(n-k) where in this case n C k = combination of n choose k n = 6 (six arrows total are shot) k = 4 (we want to know the probability of getting exactly 4) p = 0.80 (decimal form of 80%) For more info, check out http://www.stat.yale.edu/Courses/1997-98/101/binom.htm

OpenStudy (anonymous):

thanks jim

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

tell me what you get teenagerunaway

OpenStudy (anonymous):

Thank you both of you!

OpenStudy (anonymous):

i got .24

jimthompson5910 (jim_thompson5910):

same here and you can use this handy calculator to check (but it helps to know how to do it through the formula) http://stattrek.com/online-calculator/binomial.aspx

jimthompson5910 (jim_thompson5910):

well more accurately 0.24576

OpenStudy (anonymous):

how would i do a problem saying no more than

jimthompson5910 (jim_thompson5910):

"no more than" means "less than or equal to"

jimthompson5910 (jim_thompson5910):

so what you have to do is calculate the individual probabilities from k = 0 to k = max limit, then add them up

OpenStudy (anonymous):

so it would be P(x less than equal to x?

jimthompson5910 (jim_thompson5910):

more like P(x <= k) but you have the right idea

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

for instance, in that calculator I posted, it would be the line that says "Cumulative Probability: P(X ≤ 4)" (in this case, k = 4)

OpenStudy (anonymous):

Got it